OpenStudy (anonymous):

)Margie takes a random sample of 30 students in her seventh-grade class and finds that 75% of the sample prefers the color blue over the color red. There are 125 students in the seventh grade. Based on the sample proportion, about how many students in the seventh grade would be expected to prefer blue over red? 30 75 94 125

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: 75% means that every 100 students, only 75 prefer blue color over red color

OpenStudy (hitaro9):

Okay, so the 30 students is a bit of a misleading fact. It doesn't really matter for a problem like this. All you need to know is the proportion.

OpenStudy (anonymous):

so its 75?

OpenStudy (hitaro9):

So we have ~75% of students liking the color blue. Can you convert this to decimal or fractions?

OpenStudy (anonymous):

.75

OpenStudy (anonymous):

75/100

OpenStudy (hitaro9):

Good. That's going to be about the proportion of students liking the color blue. Now, we want to see, out of 125 students, about how many will like the color blue?

OpenStudy (michele_laino):

no, since we have 125 students, so we have to solve this proportion: 125: x = 100: 75 where x is the requested number what is x?

OpenStudy (anonymous):

30?

OpenStudy (amistre64):

both methods are fine

OpenStudy (hitaro9):

To do this, we want the proprortion to 125 to equal .75 so 75/100 = X/125 Do you think you can solve it from there?

OpenStudy (michele_laino):

thanks! @amistre64

OpenStudy (anonymous):

is it 30?

OpenStudy (hitaro9):

The ncie thing about multiple choice questions is you can guess and check. Why don't you try plugging 30 in for x and seeing if it equals .75?

OpenStudy (michele_laino):

in order to solve that proportion, we have to apply the fundamental property of proportions, namely: 100* x= 75*125 please divide both sides by 100 and you will get your answer

OpenStudy (anonymous):

125..? i think cuz 125/100 is 1.25

OpenStudy (michele_laino):

then multiply 1.25 by 75

OpenStudy (anonymous):

94

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thx again Michele

OpenStudy (michele_laino):

thanks! @ashley_13

OpenStudy (hitaro9):

Do you want to go over some of the reasoning again? Or do you think you could do this type of problem on your own?

OpenStudy (anonymous):

i culd do it on my own ^.^ thx